OpenStudy (anonymous):

joe math HELP!!!!!!!!! by using the first principle of differentiation differentiate y=sqrt(x+1) show all steps

OpenStudy (saifoo.khan):

HAHA-SHER!

OpenStudy (dumbcow):

\[y = \sqrt{x+1} = (x+1)^{1/2}\] \[dy/dx = (1/2)(x+1)^{(1/2-1)} = (1/2)(x+1)^{-1/2}\] \[dy/dx = \frac{1}{2\sqrt{x+1}}\]

OpenStudy (zarkon):

I believe he wants the derivative using \[\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

@dumbcrow use first principle of differentiation